OpenStudy (anonymous):

Multiply: (3x+4)(2x-4)

OpenStudy (angela210793):

=6x^2-12x+8x-16=6x^2-4x-16

OpenStudy (lalaly):

agree wit angela :D

OpenStudy (anonymous):

see, what I got was something totally different

OpenStudy (angela210793):

3x*2x+3x*(-4)+4*2x+4*(-4)=6x^2-12x+8x-16=6x^2-4x-16